# Decoupling Understanding from Reasoning via Problem Space Mapping for Small-scale Model Reasoning

### v1

@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (Grant No. 62441617), and by the Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing.
@@ Appendix A Appendix A: Proof of Theorem 1
Let $\mathcal{Q}$ be a finite set of natural language problems, viewed as distinct states $s$, and let $A$ denote the set of candidate responses. At each round $t\in\{1,\dots,T\}$, a SLM observes a problem $s_{t}\in\mathcal{Q}$, selects an action $a_{t}\in A$, and receives a reward $r(s_{t},a_{t})$. Suppose learning is performed via a state-wise Upper Confidence Bound (UCB) algorithm in a contextual bandit setting. Then, the total regret after $T$ rounds is bounded by
@@
We consider each natural language problem $x\in\mathcal{Q}$ as a distinct state $s$, and define the candidate response set as the action space $A$. At each round $t\in\{1,\dots,T\}$, the agent observes a state $s_{t}\in\mathcal{Q}$, selects an action $a_{t}\in A$, and receives a reward $r_{t}=r(s_{t},a_{t})\in[0,1]$ with expected value $\mu(s_{t},a_{t})$. Let $\mu^{*}(s)=\max_{a\in A}\mu(s,a)$ be the optimal expected reward for state $s$. The total regret is defined as:
@@
Step 2: Bound regret for a single state. Fix $s\in\mathcal{Q}$, and let $T_{s}=\sum_{a\in A}N_{T}(s,a)$ be the number of times state $s$ occurs. When in state $s$, the agent faces a standard multi-armed bandit problem with $|A|$ arms. By the UCB algorithm, the regret for state $s$ is bounded by (Auer, Cesa-Bianchi, and Fischer 2002; Bubeck, Cesa-Bianchi et al. 2012):
@@ Appendix E Appendix E: Additional Experimental Results
To further validate DURIT’s effectiveness in improving training efficiency, we compare the training rewards of DURIT (Step III of the first iteration) with those of the GRPO baseline on the Qwen2.5-0.5B-Instruct and Llama3.2-1B-Instruct models, as shown in Figures 11 and 12. For both models, DURIT achieves the same reward level as the converged GRPO baseline after only a fraction of the training steps, significantly reducing training time and improving sample efficiency. Moreover, DURIT yields a higher final reward upon convergence, demonstrating its ability not only to accelerate reinforcement learning by decoupling understanding from reasoning and compressing the effective state space, but also to improve the quality of convergence. These results also provide empirical support for Theorem 1.

### v2

@@ Acknowledgments
This work was supported by the National Science and Technology Major Project (Grant No. 2022ZD0117402), the National Natural Science Foundation of China (Grant No. 62441617), and the Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing.
@@ Appendix A Appendix A: Proof of Theorem 1
Let $\mathcal{Q}$ be a finite set of natural language problems, viewed as distinct states $s$, and let $A$ denote the set of candidate responses. At each round $t\in\{1,\dots,T\}$, a SLM observes a problem $s_{t}\in\mathcal{Q}$, selects an action $a_{t}\in A$, and receives a reward $r(s_{t},a_{t})$. Suppose learning is performed via a state-wise Upper Confidence Bound (UCB) algorithm in a contextual bandit setting. Then, in the state-independent worst case, the total regret after $T$ rounds is bounded by
@@
For each natural language problem $x\in\mathcal{Q}$, a SLM $m$ with a token limit $T$ and vocabulary $V$ has up to $V^{T}$ possible responses, forming a finite bandit problem. Thus, we treat each problem $x$ as a distinct state $s$ and define the candidate response set as the action space $A$. This formulation is also aligned with commonly adopted outcome reward theoretical analyses (Cui et al. 2025) and algorithms (Shao et al. 2024; Yu et al. 2025). At each round $t\in\{1,\dots,T\}$, the agent observes a state $s_{t}\in\mathcal{Q}$, selects an action $a_{t}\in A$, and receives a reward $r_{t}=r(s_{t},a_{t})\in[0,1]$ with expected value $\mu(s_{t},a_{t})$. Let $\mu^{*}(s)=\max_{a\in A}\mu(s,a)$ be the optimal expected reward for state $s$. The total regret is defined as:
@@
Step 2: Bound regret for a single state. Fix $s\in\mathcal{Q}$, and let $T_{s}=\sum_{a\in A}N_{T}(s,a)$ be the number of times state $s$ occurs. When in state $s$, the agent faces a standard multi-armed bandit problem with $|A|$ arms. By the UCB algorithm, in the state-independent worst case, the regret for state $s$ is bounded by (Auer, Cesa-Bianchi, and Fischer 2002; Lattimore and Szepesvári 2020):
@@ Appendix E Appendix E: Additional Experimental Results
Appendix E.1: Additional Iterative Experiments
To investigate the convergence behavior and effectiveness of DURIT’s iterative training, we conducted experiments on the GSM8K dataset using the Qwen2.5-0.5B-Instruct and Llama3.2-1B-Instruct models. As shown in Table 6, the average accuracy of both models generally improves with the number of iterations in most cases, confirming the benefits of iterative refinement; however, performance gains gradually diminish due to saturation, typically converging after approximately three iterations, while iterative training on different datasets still exhibits potential for further improvement. To further validate this effect under equal training budgets, we compared two strategies on the Qwen2.5-0.5B-Instruct base model: performing two iterations versus a single iteration with double the training steps (DTT). The results show that, under the same total training time, a single DTT iteration achieves lower accuracy than two iterations, further corroborating the advantage of DURIT’s iterative approach.
Methods In-Domain Out-of-Domain Average
GSM8K-Platinum MAWPS SVAMP MATH500 GAOKAO LogiQA
# Qwen2.5-0.5B-Instruct based
DURIT (iter=1) 53.68 60.19 62.67 31.00 23.39 24.58 42.59
DURIT (iter=1, DTT) 54.43 60.58 60.00 31.40 24.19 25.19 42.63
DURIT (iter=2) 53.10 60.38 63.00 32.80 22.58 25.81 42.95
DURIT (iter=3) 53.76 60.58 63.67 31.00 24.19 25.65 43.14
DURIT (iter=4) 54.34 61.15 64.33 31.00 21.77 26.27 43.14
# Llama3.2-1B-Instruct based
DURIT (iter=1) 50.37 59.62 64.33 26.00 14.52 21.20 39.34
DURIT (iter=2) 52.36 62.31 66.00 27.60 12.10 19.82 40.03
DURIT (iter=3) 52.85 61.54 68.00 27.80 13.71 15.67 39.93
DURIT (iter=4) 53.43 61.76 67.00 26.60 14.52 17.05 40.06
Table 6: Iterative performance (%) of DURIT on six benchmarks using Qwen2.5-0.5B-Instruct and Llama3.2-1B-Instruct as base models. Bold and underline indicate the best and second-best results, respectively. DTT denotes double training time for a single iteration.
@@
We also analyzed the sensitivity of $\alpha_{1}$ and $\alpha_{2}$. As shown in Figures 11 and 12, across a threefold variation in each parameter, the model’s accuracy fluctuates by no more than 0.5%, demonstrating that DURIT exhibits substantial robustness to reasonable parameter choices. The best performance is achieved when $\alpha_{1}$=1e-3 and $\alpha_{2}$=1e-2. Accordingly, we adopt these values as the default settings.
Figure 11: Average accuracy of Qwen2.5-0.5B-Instruct across six datasets under different distillation coefficient $\alpha_{1}$.
Figure 12: Average accuracy of Qwen2.5-0.5B-Instruct across six datasets under different distillation coefficient $\alpha_{2}$.
@@
To further validate DURIT’s effectiveness in improving training efficiency, we compare the training rewards of DURIT (Step III of the first iteration) with those of the GRPO baseline on the Qwen2.5-0.5B-Instruct and Llama3.2-1B-Instruct models, as shown in Figures 13 and 14. For both models, DURIT achieves the same reward level as the converged GRPO baseline after only a fraction of the training steps, significantly reducing training time and improving sample efficiency. Moreover, DURIT yields a higher final reward upon convergence, demonstrating its ability not only to accelerate reinforcement learning by decoupling understanding from reasoning and compressing the effective state space, but also to improve the quality of convergence. These results also provide empirical support for Theorem 1.